# Machine Learning, Density Functional Theory, and Experiments to Understand the Photocatalytic Reduction of CO by CuPt/TiO

### 14 Feb 2024 v1

@@ Abstract
Despite the challenges of low-yield ( mol) for large-scale applications in photocatalytic CO conversion to hydrocarbons, the activity and selectivity of TiO-based systems [14, 15, 13, 16] are encouraging. Since pristine TiO suffers from a low yield due to several reasons, including the maximum-visible light under-utilization, [17] large electron-hole separation, [10] and wide band gaps 3.0 eV (rutile)[18] and 3.2 eV (anatase),[19] several strategies have been tested insofar. Manipulating the geometric/electronic properties of TiO via creation of oxygen vacancies or depositing metal (e.g. Cu/Pt) nanoparticles [20] have potential to enhance the photocatalytic activity.
On TiO surfaces, CO activation is a bottleneck step owing to a variety of reasons. [21, 22, 23, 24] For instance, the *CO formation via absorption of photo-excited electrons: CO *CO+e is thermodynamically hindered whilst the *CHOO formation via absorption of two electrons: CO+H *CHOO+2e is kinetically hindered.[24] The trivial charge transfer between *CO and TiO would result in weaker interaction.[21, 22, 23] Since CO weakly interacts with defect-laden TiO (see SI, section S1) it cannot facilitate CO activation without external driving force, thus necessitating an engineering of TiO structures.
@@
For the accelerated discovery of materials for CO reduction, various machine learning (ML) models have been utilized to study several systems [31, 32, 33, 34] including (bare) semiconductor oxides.[32] More recently, E(3)-equivariant neural network (ENN) is attracted attention owing to high data efficiency for ML models.[35, 36] The ML approach in combination with basin-hopping Monte Carlo (BHMC) method can provide an efficient way to tackle the combinatorial problem of composition of bimetallic Cu-Pt nanoclusters together with CO adsorbates, relating to the large degrees of freedom due to the number of adsorption sites and the orientation of clusters and CO.
Whilst the ab-initio methods offer atomistic insights into the potential energy surface of material systems but suffer from the high computational cost and inefficient scaling with system sizes, ML interatomic potentials (MLIPs) can provide substantial speed-ups at the expense of accuracy with linear scaling. ML can often reduce computational costs by orders of magnitude.[37, 38] MLIPs have been improved drastically to achieve higher accuracy in energies ( meV/atom) and forces (0.1 eV/Å) and have been utilized in several applications, e.g. in understanding complex materials chemistry, [39, 40, 41, 42, 43, 44], enumerating nanostructures configurations,[41, 45] performing long-time MD simulations,[46, 47, 48] and exploring catalytic reactions.[49, 50, 51, 52] Nevertheless, MLIPs suffer from highly inhomogeneous feature-space sampling of training set (inherent biases). The inclusion of underrepresented configurations leads to significant errors since the data representing inhomogeneous local environments are overshadowed by core atoms.[53] To model CO@CuPt/TiO systems, it thus requires an inherent-bias correction.
@@ 2.2 Potential energy surface exploration for CO adsorption on CuPt/TiO
The extremely large-configurational space is apparent from an illustration of PES exploration for a model system (Fig.2), whose configurations are found to be different cluster shapes with various arrangement of Pt/Cu atoms together with several possible adsorption sites for CO, including the CuPt/TiO interfacial sites as well as several different oxygen vacancy sites on TiO(110).
@@ 2.3 Interaction of CuPt nanocluster with TiO
For CuPt composition, the cluster-support interaction is facilitated mostly by chemical bonds formed between Cu and bridging O atoms and bonds between Pt and bridging O and/or five-fold () coordinated Ti atoms. The distributions of Cu-O bond lengths for the lowest energy configurations (Fig.S4) and more details can be found in SI (see S2.4). Owing to the metal-support interaction, the CuPt cluster adopts a hemi-spherical shape (Fig.3). Regardless of the composition, the supported CuPt cluster prefers a 3D geometrical shape.
It is important to understand the distribution of CuPt/TiO frontier states since the states spatially distributed around interfacial atoms would play an important role in ultrafast dynamics and thereby photo-induced reactions.[54] For CuPt/TiO (), the frontier states are spattially localized at interface and around Cu-Pt atoms (Fig.3). Although the spatial localization is found to vary with Cu-Pt composition, the common feature is that there is some distribution of maximum VB frontier states. In contrast, Cu/TiO frontier states localize at the TiO second layer (from the top). These results suggest that the CuPt/TiO interface would be more active than Cu/TiO interface.
@@ 2.4 CO interaction with CuPt/TiO
We now examine the interaction of CO with CuPt/TiO via DFT by considering the global minima configurations predicted by MLIP+BHMC simulations. For all CuPt compositions, we find that CO adsorption is thermodynamically favorable, and CO is activated (Table 1). Regardless of the CuPt composition, CO adsorbs at the interface and the degree of interaction depends upon the composition and shape of clusters. The CO interaction with Pt/TiO is stronger as compared to Cu/TiO systems. Although we do not observe any specific trend in adsorption energies, our calculations suggest that the CO interaction magnitude would be higher for CuPt clusters than for only pure Cu. As compared to the bond length (1.162 Å) and angle (180) of CO in gas phase, upon its adsorption the C-O bond elongates, and the bond angle changes to 121-134. Taken together results of the CO interaction with CuPt/TiO systems with various Pt/Cu compositions (=-0.50 to -1.2 eV), the charge transfer to an adsorbed CO (0.7 to 0.9) and the elongation of C-O bond (0.2 Å) and the substantial reduction of OCO (55-59), we show that CuPt/TiO systems facilitate CO activation.
@@
Fig.4 and others in Fig.S5 (SI, S2.4) present the global-minima structures of CO@CuPt/TiO configurations. CO adsorbs at the interface of Cu/TiO such that C attaches to a Cu atom and O atom fills out the O vacancy site (Fig.4a). For CuPt with (Pt 1), the *CO adsorbs at the interface with an O atom occupying O vacancy site and C bonded to Pt, thus preferring a Pt over Cu site (Fig.4b-e). For , CO adsorbs in a way that C is bonded to Pt and a O to Ti (Fig.4f). The CO adsorption geometries in the case of are like that of . These results indicate that the CuPt/TiO interface serves as active sites for CO adsorption, particularly highlighting the importance of interfacial Cu/Pt and surface Ti atoms (near O vacancy).
From the electron density difference (n) plot (Fig.4g), it is seen that the resulting electron density is re-distributed in the bonding regions of CO atoms (C, O1, O2) and interfacial atoms (Pt/Ti). Thus, the favorable interaction between CO and CuPt/TiO is mediated by the interface. The charge re-distribution, upon CO interaction with CuPt/TiO, suggests that the charge transfer takes place between two sub-systems. There is charge accumulation around O atoms and charge depletion around a C atom and surface Ti atoms. Some charge addition to the *CO anti-bonding state can trigger the C-O bond elongation. Indeed, based on the partial charges analysis, the net charge transfer of 0.96e takes place from CuPt/TiO to an adsorbed CO (Table 1), inducing the bond elongation and thereby facilitating CO activation.
Hereinafter, we present the results for structure and energetics properties of CuPt/TiO model system when interacting with the reactants (CO/HO) and key intermediates of CO reduction. The model selection is partly based on the stronger CO adsorption at interfacial sites of CuPt/TiO and partly on the higher Pt than Cu (prevent a CuPt cluster from oxidizing).
@@ 2.5 Energetics for CO dissociation vs hydrogenation
Based on the DFT-calculated energetics, the *CO formation from *CO dissociation is energetically favorable (E=-0.62 eV). In contrast, the CO hydrogenation via *CO+*H *CHOO+* step is endothermic (E=+0.60 eV). Similarly, the *CO+*H*COOH+* is endothermic (E=+0.27 eV). Although these steps require some energy, the reactions might be activated in normal reaction conditions. Calculated energetics provide some indication that while *CO formation is a favorable process (*CO hydrogenation can proceed to yield *CHO formation, being more thermodynamically favorable than *COH formation.[55]), the *CO hydrogenation is also feasible under external driving force.

### 16 Feb 2024 v2

@@ Abstract
Despite the challenges of low-yield ( mol) for large-scale applications in photocatalytic CO conversion to hydrocarbons, the activity and selectivity of TiO-based systems [14, 15, 13, 16] are encouraging. Since pristine TiO suffers from a low yield due to several reasons, including the maximum-visible light under-utilization,[17] large electron-hole separation,[10] and wide band gaps 3.0 eV (rutile)[18] and 3.2 eV (anatase),[19] several strategies have been tested insofar. Manipulating the geometric/electronic properties of TiO via creation of oxygen vacancies or depositing metal (e.g. Cu/Pt) nanoparticles have potential to enhance the photocatalytic activity.[20]
On TiO surfaces, CO activation is a bottleneck step owing to a variety of reasons.[21, 22, 23, 24] For instance, the *CO formation via absorption of photo-excited electrons: CO *CO+e is thermodynamically hindered whilst the *CHOO formation via absorption of two electrons: CO+H *CHOO+2e is kinetically hindered.[24] The trivial charge transfer between *CO and TiO would result in weaker interaction.[21, 22, 23] Since CO weakly interacts with defect-laden TiO (see section S1 and Fig.S1 in supporting information) it cannot facilitate CO activation without external driving force, thus necessitating an engineering of TiO structures.
@@
For the accelerated discovery of materials for CO reduction, various machine learning (ML) models have been utilized to study several systems [31, 32, 33, 34] including (bare) semiconductor oxides.[32] More recently, E(3)-equivariant neural network (ENN) is attracted attention owing to high data efficiency for ML models.[35, 36] The ML approach in combination with basin-hopping Monte Carlo (BHMC) method can provide an efficient way to tackle the combinatorial problem of composition of bimetallic Cu-Pt nanoclusters together with CO adsorbates, relating to the large degrees of freedom due to several adsorption sites and the orientation of clusters and CO.
Whilst the ab-initio methods offer atomistic insights into the potential energy surface (PES) of material systems but suffer from the high computational cost and inefficient scaling with system sizes, ML interatomic potentials (MLIPs) can provide substantial speed-ups at the expense of accuracy with linear scaling. ML can often reduce computational costs by orders of magnitude.[37, 38] MLIPs have been improved drastically to achieve higher accuracy in energies ( meV/atom) and forces (0.1 eV/Å) and have been utilized in several applications, e.g. in understanding complex materials chemistry,[39, 40, 41, 42, 43, 44], enumerating nanostructures configurations,[41, 45] performing long-time MD simulations,[46, 47, 48] and exploring catalytic reactions.[49, 50, 51, 52] Nevertheless, MLIPs suffer from highly inhomogeneous feature-space sampling of training set (inherent biases). The inclusion of underrepresented configurations leads to significant errors since the data representing inhomogeneous local environments are overshadowed by core atoms.[53] To model CO@CuPt/TiO systems, it thus requires an inherent-bias correction.
@@ 2.2 Potential energy surface exploration for CO adsorption on CuPt/TiO
The extremely large-configurational space is apparent from an illustration of PES exploration for a model system (Fig.2), whose configurations are found to be different cluster shapes with various arrangement of Pt/Cu atoms together with several possible adsorption sites for CO, including the CuPt/TiO interfacial sites as well as different oxygen vacancy sites on TiO(110).
@@ 2.3 Interaction of CuPt nanocluster with TiO
For CuPt composition, the cluster-support interaction is facilitated mostly by chemical bonds formed between Cu and bridging O atoms and bonds between Pt and bridging O and/or five-fold () coordinated Ti atoms. The distributions of Cu-O bond lengths for the lowest-energy configurations (Fig.S4) and more details can be found in SI (see S2.3). Owing to the metal-support interaction, the CuPt cluster adopts a hemi-spherical shape (Fig.3). Regardless of the composition, the supported CuPt cluster prefers a 3D geometrical shape.
It is important to understand the distribution of CuPt/TiO frontier states since the states spatially distributed around interfacial atoms would play an important role in ultrafast dynamics and thereby photo-induced reactions.[54] For CuPt/TiO (), the frontier states are spattially localized at interface and around Cu-Pt atoms (Fig.3). Although the spatial localization is found to vary with Cu-Pt composition, the common feature is that there is some distribution of maximum valence band (VB) frontier states. In contrast, Cu/TiO VB frontier states localize at the TiO second layer (from the top). These results suggest that the CuPt/TiO interface would be more active than Cu/TiO interface.
@@ 2.4 CO interaction with CuPt/TiO
We now examine the interaction of CO with CuPt/TiO via DFT by considering the lowest-minima configurations predicted by MLIP+BHMC simulations. For all CuPt compositions, we find that CO adsorption is thermodynamically favorable, and CO is activated (Table 1). Regardless of the CuPt composition, CO adsorbs at the interface and the degree of interaction depends upon the composition and shape of clusters. The CO interaction with Pt/TiO is stronger as compared to Cu/TiO systems. Although we do not observe any specific trend in adsorption energies, our calculations suggest that the CO interaction magnitude would be higher for CuPt clusters than for only pure Cu. As compared to the bond length (1.162 Å) and angle (180) of CO in gas phase, upon its adsorption the C-O bond elongates, and the bond angle changes to 121-134. Taken together results of the CO interaction with CuPt/TiO systems with various Pt/Cu compositions (=-0.50 to -1.2 eV), the charge transfer to an adsorbed CO (0.7 to 0.9) and the elongation of C-O bond (0.2 Å) and the substantial reduction of OCO (55-59), we show that CuPt/TiO systems facilitate CO activation.
@@
Fig.4 and others in Fig.S5 (SI, S2.4) present the lowest-minima structures of CO@CuPt/TiO configurations. CO adsorbs at the interface of Cu/TiO such that C attaches to a Cu atom and O atom fills out the O vacancy site (Fig.4a). For CuPt with (Pt 1), the CO adsorbs at the interface with an O atom occupying O vacancy site and C bonded to Pt, thus preferring a Pt over Cu site (Fig.4b-e). For , CO adsorbs in a way that C is bonded to Pt and a O to Ti (Fig.4f). The CO adsorption geometries in the case of are like that of . These results indicate that the CuPt/TiO interface serves as active sites for CO adsorption, particularly highlighting the importance of interfacial Cu/Pt and surface Ti atoms (near O vacancy).
From the electron density difference () plot (Fig.4g), it is seen that the resulting electron density is re-distributed in the bonding regions of *CO atoms (C, O1, O2) and interfacial atoms (Pt/Ti). Thus, the favorable interaction between CO and CuPt/TiO is mediated by the interface. The charge re-distribution, upon CO interaction with CuPt/TiO, suggests that the charge transfer takes place between two sub-systems. There is charge accumulation around O atoms and charge depletion around a C atom and surface Ti atoms. Some charge addition to the *CO anti-bonding state can trigger the C-O bond elongation. Indeed, based on the partial charges analysis, the net charge transfer of 0.96e takes place from CuPt/TiO to an adsorbed CO (Table 1), inducing the C-O bond elongation and thereby facilitating CO activation.
Hereinafter, we present the results for structure and energetics properties of CuPt/TiO model system when interacting with the reactants (CO/HO) and key intermediates of CO reduction. The model selection is partly based on the stronger CO adsorption at interfacial sites of CuPt/TiO and partly on the higher Pt concentration than Cu (prevent a CuPt cluster from oxidizing).
@@ 2.5 Energetics for CO dissociation vs hydrogenation
Based on the DFT-calculated energetics, the *CO formation from *CO dissociation is energetically favorable (E=-0.62 eV). In contrast, the CO hydrogenation via *CO+*H *CHOO+* step is endothermic (E=+0.60 eV). Similarly, the *CO+*H*COOH+* is endothermic (E=+0.27 eV). Although these steps require some energy, the reactions might be activated in normal reaction conditions. Calculated energetics provide some indication that while *CO formation is a favorable process (*CO hydrogenation can proceed to yield *CHO formation, being more thermodynamically favorable than *COH formation.[55]), the *CO hydrogenation might also feasible under external driving force.